### v3

# Spectral Mode Enhancement in Coherent-harmonic Dual-comb Spectroscopy Enables Exceeding 300-fold averaging-time reduction

Abstract
Dual-comb spectroscopy (DCS) is a novel Fourier-transform spectroscopy not relying on mechanical scanning and capable of simultaneously achieving high speed, high spectral resolution, and broad optical bandwidth. Nevertheless, it suffers from low signal-to-noise ratio (SNR) per single acquisition due to the dynamic range limitation of photodetectors imposed by the high-peak-power mode-locked pulses, making coherent averaging an essential means to improve SNR, at the price of compromising the exceptional time resolution and placing more stringent demands on mutual coherence and stability. In this study, a novel approach to enhance SNR by exploiting the spectral mode enhancement mechanism in coherent-harmonic pulses is demonstrated. As a proof-of-concept, two frequency combs with mode spacing of $\sim$12.5 MHz, operated at a 20th harmonic repetition rate of $\sim$250 MHz, are employed, demonstrating a $>$300-fold reduction in averaging time for comparable SNR in conventional DCS. This reduction is expected to be further enhanced through integration with ultra-high repetition rate combs like microresonator combs. This new approach promises both a recovery of the inherent high-speed capability and a mitigation of the coherence-time requirements, thereby making it possible to significantly facilitate subsequent DCS investigations and field deployments.
I Introduction
Dual-comb spectroscopy (DCS) is a remarkable novel Fourier-transform spectroscopy technique powered by optical frequency comb (OFC) 1. Over the past decade, it has been increasingly recognized as an important spectroscopic tool in fields that require high speed, broad bandwidth, high frequency resolution, and high frequency accuracy, such as in investigations of physical turbulent environments 2, 3, 4, 5, 6 and chemical and biological dynamic processes 7, 8, 9, 10. By employing two coherent OFCs with a slight repetition rate difference ($\Delta f_{\mathrm{rep}}$), and inherently free from mechanical moving parts, DCS can capture an entire interferogram (IGM), yielding a spectrum with a full spectral resolution set by the repetition rate of the combs. The time frame for this capture is as short as $1/\Delta f_{\mathrm{rep}}$, and with large $\Delta f_{\mathrm{rep}}$ configurations, this leads to theoretical temporal resolution reaching microsecond or even nanosecond timescales 11, 12, 13.
However, in typical practical DCS systems, the high peak power of mode-locked pulses often leads to photodetector saturation or nonlinear response, thus restricting the average power of the combs to a low level. Consequently, given the vast number of comb teeth inherent to broadband spectra, the power per comb tooth is typically limited to the nanowatt range 14, 15, ultimately resulting in extremely low SNR per single acquisition, which necessitates coherent averaging or multi-period acquisition to enhance the SNR 16, 15, 17, 12. Unfortunately, the SNR gain achieved through averaging scales proportionally to the square root of the number of averages. For instance, to improve the SNR by a factor of 10, a hundred averages are needed. Obviously, this process not only significantly compromises the exceptional time-resolution of DCS by several orders of magnitude, but also places greater demands on the long-term mutual coherence and stability of the system, which greatly increases the implementation complexity and cost.
To mitigate this issue, one approach involves employing sequential spectrum acquisition using multiple photodetectors 14 or averaging spectral copies instead of IGMs 18. However, the SNR enhancement offered by these two methods has been limited by experimental complexity or sampling bandwidth constraints, demonstrating averaging-time reduction of less than one order of magnitude. Another early approach is to stretch the pulses of the probe comb using a chirped fiber Bragg grating 19. This method can effectively reduce the dynamic range of an IGM; however, it has limited impact on the nonlinearity or saturation induced by pulses, as the detector’s response time significantly exceeds the pulse duration. More recently, the utilization of $\sim$ 1-GHz-repetition-rate combs 12, 20, 13, 21 offers higher average power and larger $\Delta f_{\mathrm{rep}}$. Nevertheless, a $\sim$ 1-GHz repetition rate had struck a delicate balance between measurement speed and spectral resolution 22. A higher repetition rate inherently implies a lower spectral resolution, which limits further increases in repetition rate.
From this perspective, employing frequency combs analogous to certain harmonic mode-locked lasers 23, whose pulse repetition rate ($f_{\mathrm{rep}}$) operates on the harmonics of the laser cavity’s fundamental repetition rate, thus possibly featuring a high $f_{\mathrm{rep}}$ but a low mode spacing ($f_{\mathrm{ms}}$), may offer a promising avenue. On the one hand, this configuration can enable higher average power within the same dynamic range. On the other hand, the phenomenon that the pulse repetition rate exceeding the mode spacing likely points to an underlying mechanism of the harmonic pulses coherently enhancing their common spectral modes. If this mechanism can be effectively transferred to interferograms in DCS, i.e., generate multiple centerbursts within the time period of $1/\Delta f_{rep}$ to coherently enhance RF comb modes, it would be possible to achieve the desired SNR gain, thereby potentially refraining from extended IGMs coherent averaging.
Here we find pulse temporal phase modulation (TPM) using electro-optical phase modulator (EOPM) provides an ideal means to realize such combs unaffected by supermode noise 32. This TPM has been employed for the comb densification technique 25, 26, 27, 28, 29, 30, 31 based on spectral self-imaging effect 24 over the past decade or so, which involves applying periodic TPM to the pulses, inducing evenly spaced ”self images” of the original comb lines on the spectrum, thereby densifying the comb. In 2016, N.B. Hébert et al. 33 first introduced this densification technique in DCS and presented an enhanced sub-MHz spectral resolution from a 100-MHz comb, using a pseudo-random binary sequence (PRBS) modulation. In 2024, C. Q.-Galan et al. 34 extended this method to 1-GHz gain-switched OFCs, achieving $\sim$ 3.9 MHz resolution. Despite this, both of these two studies primarily focused on the spectral resolution enhancement, merely treating such combs as conventional OFCs with small mode spacing. The underlying harmonic-enhancement mechanism and the descriptions of such special dual-comb have not been elucidated.
In this work, we derived a coherent-harmonic-pulse superposition model in the frequency domain to elucidate this mechanism, and presented the time-domain and equivalent frequency-domain descriptions of two architectures for this special DCS, with a focus on their respective SNR gains. To align with its intuitive physical picture and distinguish it from conventional dual-comb spectroscopy (C-DCS), we term this approach coherent-harmonic-enhanced dual-comb spectroscopy (CH-EDCS). As a proof of concept, we perform transmission spectrum measurements of a hydrogen cyanide ($\mathrm{H^{13}C^{14}N}$) gas cell the two proposed architectures and, for comparison, a C-DCS architecture. The experiments were based on two commercial 250 MHz combs, which were configured using TPM and a pulse-picking technique 41 to operate at a spectral resolution of 12.5 MHz. As a result, we achieved an SNR gain of up to $\sim$ 18 times per single acquisition, equivalent to a $>$300-fold reduction in averaging time, demonstrating the significant potential for SNR enhancement of this approach.
Fig. 1: Concepts of conventional DCS (C-DCS) and C-H-DCS. (a) Time-domain description of C-DCS. The pulse train (red) of comb 1 progressively steps across another pulse train (blue) with a slight repetition rate difference $\Delta f_{\mathrm{s}}$, generating an interferogram with a single centerburst in a period of $1/\Delta f_{\mathrm{s}}$. When the temporal interval between the two pulses is large, effective interference does not occur.(b) The equivalent frequency-domain description of the C-DCS. The beating of two combs produces an RF comb in the radio-frequency domain. (c) Time-domain description of the C-H-DCS. The pulse of the conventional optical frequency comb (C-OFC) with a pulse repetition rate of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$, steps across the four pulses of the coherent-harmonic optical frequency comb (H-OFC), which have phases of $[\pi,0,0,0]$ and a pulse repetition rate of $4f_{\mathrm{ms}}$. This generates four centerbursts within an interferogram period of $1/\Delta f_{\mathrm{s}}$, which inherit the $[\pi,0,0,0]$ phases. (d) The equivalent frequency-domain description of the C-H-DCS. Two combs with mode spacing of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$ and $f_{\mathrm{ms}}$ respectively beat together, producing an RF comb in the radio-frequency domain. This RF comb inherits the $[0,-\pi,-\pi,-\pi]$ periodic multilevel phase profile of the H-OFC and exhibits twice the line amplitude compared to the C-DCS.
II Principles
In this section, two DCS architectures for this CH-EDCS concept are studied: one termed C-H-OFC, employing a coherent-harmonic optical frequency comb (H-OFC) realized by TPM and a conventional optical frequency comb (C-OFC); the other termed H-H-DCS, employing two H-OFCs. For clarity, we use a harmonic number of m = 4 and a TPM sequence of $[\pi,0,0,0]$ as illustrative examples.
II.1 C-H-DCS
As illustrated in Figs. 1 (c) and (d), C-H-DCS utilizes a H-OFC whose repetition rate is approximately equal to 4 times of the repetition rate of the C-OFC (or, more generally, any integer multiple m). The pulses of comb 1, after stepping across the first pulse of the H-OFC, subsequently scan its 2nd to 4th pulses, thus generating four centerbursts instead of a single one on the interferogram within the same time period. Additionally, according to the principle of linear optical sampling 35, 36, the expression for the interference intensity can be written as
$I(n)\sim\sum_{n}\langle|E_{1}(t-n\tau)||E_{2}(t)\left|\cos\left(\Delta\omega_{c}t+\phi_{TPM}\right)\right\rangle,$ (1)
where $E_{1}(t),E_{2}(t)$ represents the pulses’ electric field envelopes; n denotes the pulse pair number; $\tau$ denotes the time delay induced by the repetition rate difference; and $\phi_{TPM}$ represents the phase difference introduced by the TPM. For each centerburst, the $\phi_{TPM}$ is constant. Consequently, the TPM imposed on the pulses of H-OFC is finally mapped onto the four centerbursts.
Fig. 2: Illustration of the process of the spectral mode enhancement by the harmonic pulse trains in the frequency domain. The j represents the spectral mode index (derived from $\omega_{j}=j\cdot 2\pi f_{\mathrm{ms}}$). The colored linear arrows with indexes of k = 1, 2, 3, 4 represent the mode vectors from the kth-sub-harmonic pulse train at the frequency of $\omega_{j}$, while the blue rotating arrows indicate the temporal phase modulation (TPM), which rotates all the modes of the 1st-sub-harmonic pulse train (k = 1) by $\pi$. In the absence of this TPM, only the modes of $j=4n$ $(n\in\mathbb{Z}^{+})$ will be present, while the remaining modes will undergo destructive interference, consistent with conventional expectation. The black arrows represent the total mode vectors after superposition, which have twice the amplitude of the sub-modes and exhibit a periodic phase profile of $[-\pi,-\pi,-\pi,0]$.
Fig. 3: Concept of H-H-DCS. (a) Time-domain description of the H-H-DCS. The H-OFC 1 and H-OFC 2 emit two pulse trains with repetition rate of $4(f_{\mathrm{ms}}-\Delta f_{\mathrm{s}})$ and $4f_{\mathrm{ms}}$ respectively, both with pulse phases of $[\pi,0,0,0]$. Interference of the two pulse trains within a time period of $1/\Delta f_{\mathrm{s}}$ produces four distinct centerbursts: one at the fundamental oscillation frequency (purple), exhibiting a 4-fold amplitude enhancement over conventional dual-comb spectroscopy (C-DCS); and three additional centerbursts at higher-order frequencies (pink), which are suppressed by low-pass filtering in practical implementations. (b) The equivalent frequency-domain description of the H-H-DCS. Two H-OFCs with mode spacing of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$ and $f_{\mathrm{ms}}$ respectively beat together, producing an RF comb in the radio-frequency domain which has a uniform phase profile and exhibits four times the line amplitude compared to the C-DCS. Note that, the uniform phase profile requires the phase profiles of the two H-OFCs to be accurately aligned.
The spectral behavior of the H-OFC and the interferogram with 4 centerbursts can be elucidated through the interference of the harmonic pulses. The original pulse train is decomposed into a collection of four sub-harmonic pulse trains, each with a period of $1/f_{\mathrm{ms}}$ ($1/\Delta f_{\mathrm{s}}$ for the interferogram) and corresponding to a sub-spectrum with mode spacing of $f_{\mathrm{ms}}$. According to the linear property of the Fourier transform, the spectrum of the H-OFC or the interferogram at frequency $\omega_{j}$ is a vector superposition of the 4 sub-modes at that frequency, given by $E_{j}=\sum_{k=1}^{4}a_{0}\exp{(\omega_{j}t+\phi_{k,j})}$, where $a_{0}$ is the amplitude of the sub-modes. Considering the influence of the relative time delay among these harmonic trains and the TPM, the phase spectrum of the kth-sub-spectrum can be expressed as
$\phi_{k,j}=-\omega_{j}t_{k}+\phi_{TPM,k}=-\frac{j}{4}(k-1)\cdot 2\pi+\phi_{TPM,k}.$ (2)
The first term in the formula, $-\omega_{j}t_{k}$, arises from the contribution of the Fourier time shift, where $\omega_{j}=j\cdot 2\pi f_{\mathrm{ms}}$ is the angular frequency of the $j_{th}$ mode, $t_{k}=(k-1)/4f_{\mathrm{ms}}$ is the time delay of kth-sub-harmonic pulse train relative to the first sub-harmonic pulse train; the second term, $\phi_{TPM,k}$, is the phase shift added to the kth-train through the TPM. The superposition process and results are illustrated in Fig. 2, with vector sum $E_{j}=2a_{0}\exp{(\omega_{j}t+\varphi_{j})}$ where $\varphi_{j}$ is a periodic sequence $[-\pi,-\pi,-\pi,0]$. This indicates that the spectrum of the original pulse train has the same mode spacing but twice the amplitude of the sub-spectrum, and exhibits a periodic multilevel phase spectrum of $[-\pi,-\pi,-\pi,0]$.
Thus, an SNR gain of 2 can be achieved without coherent averaging 4 IGMs (spanning a total time of $4/\Delta f_{\mathrm{s}}$ in the C-DCS). For a more general case, the number 4 in the aforementioned equations corresponds to m, and the corresponding SNR gain is $\sqrt{m}$. Moreover, it is worth noting that if a flat gain is required, an equation in the quadratic function form indicated by the so-called ”Talbot condition” 37 can be used, which is expressed as
$\phi_{n;s,m}=\pi\frac{s}{m}k^{2}(mod~{}2\pi),$ (3)
where s, m are mutually prime integer numbers, and k is the pulse number. Correspondingly, the periodic multilevel phase profile of the total spectrum is $\varphi_{k;p,q}=-\pi\frac{p}{q}j^{2}(mod~{}2\pi),$ where $q=m$, j is the mode number, and $p\neq s$ in general. Whereas, if a reference channel for spectral normalization in DCS configuration is available, the majority of periodic signal can be set.
II.2 H-H-DCS
As illustrated in Fig. 3, C-H-DCS utilizes two H-OFCs with pulse repetition rates of $4f_{\mathrm{ms}}$ and $4(f_{\mathrm{ms}}-\Delta f_{\mathrm{s}})$. In the time domain, similar to a C-DCS with the same repetition rate difference of $4\Delta f_{\mathrm{s}}$, this architecture generates 4 centerbursts within a time period of $1/\Delta f_{\mathrm{s}}$. Due to this repetition rate difference, the phase-inverted pulses in the two pulse trains gradually slip relative to each other, resulting in distinct in-pulse-pair phase configurations: $[\pi,0,0,0]-[\pi,0,0,0]$, $[0,\pi,0,0]-[\pi,0,0,0]$, $[0,0,\pi,0]-[\pi,0,0,0]$, and $[0,0,0,\pi]-[\pi,0,0,0]$, respectively, in the time domain regions corresponding to the four centerbursts. Consequently, the $\phi_{TPM}$ in Eq. 1, is constantly 0 for the first centerburst; while for the other three centerbursts, the values are periodic $[\pi,\pi,0,0],[\pi,0,\pi,0]$, and $[\pi,0,0,\pi]$, respectively. These time-varying phases (related to the pulse pair index, i.e., time) increase the oscillation frequency of the interference electric field, shifting the 2nd to 4th centerbursts to higher-order components, which can be filtered out by an electronic low-pass filter in practical implementation. In fact, these high-order components correspond to spectral copies generated by the heterodyning of teeth from comb 1 with non-nearest teeth from comb 2 in the spectrum, as detailed in Supplement 1.
Thus, the IGM exhibits only one fundamental centerburst within a time period of $1/\Delta f_{\mathrm{s}}$, similar to that in the C-DCS. Distinctly, the fundamental centerburst originates from the contribution of 4 times the number of pulse pairs, whose interference intensity is given by $I^{\prime}(n)\sim\sum_{n=1}^{4n_{0}}\langle|E_{1}(t-n\tau/4)||E_{2}(t)|\cos\Delta\omega_{c}t\rangle$, where $n_{0}$ is the number of pulse pairs contributing to the centerburst in the C-DCS illustrated in Fig. 1(a). Consequently, following convolution with the detector response and low-pass filtering, the amplitude of this centerburst is also enhanced by a factor of 4.
In the frequency domain, two H-OFCs with 2-fold mode amplitude enhancement and periodic multilevel phase spectra of $[-\pi,-\pi,-\pi,0]$, beat together to yield an RF comb. Since the amplitude of a heterodyne beat note is the product of the amplitudes of the two beating components, and the phase is the phase difference between them, the modes of this RF comb therefore exhibit a 4-fold enhancement in amplitude and possess a uniform phase, which precisely corresponds to the single fundamental centerburst with a 4-fold increase in amplitude in the time domain.
Thus, this architecture leads to a 4-fold SNR gain. For a more general case, this number 4 corresponds to m, and is equivalent to an SNR gain brought by coherent averaging spanning a total time of $m^{2}$-fold.
Interestingly, the aforementioned IGM characteristics can be changed (the SNR gain is not changed) when the two TPM signals use the quadratic form described in Eq. 3 with the same m but different s values. The generated RF comb can exhibit quadratic periodic multilevel phase profiles of $-\pi\frac{p_{1}-p_{2}}{m}j^{2}$. According to the temporal self-imaging effect 38, 30, the fundamental centerburst on the IGM will then be redistributed, similar to that in C-H-DCS.
III Experiment
Fig. 4: Sketch of experimental setup. Rb clock: rubidium clock; OFC: optical frequency comb; AWG: arbitrary waveform generator; RFA: RF amplifiers; EOPM/EOIM: electro-optical phase/intensity modulator; OC: optical coupler; PD: photodetector; LPF: low-pass filter.
Fig. 5: Comparison of interferograms (IGMs) and signal-to-noise ratio (SNR) for the three architectures. (a) IGMs of the C-DCS and (b) its spectrum (partial). The centerburst repetition rate is 1 kHz, matching the spectral line spacing, as expected for conventional DCS. The spectrum exhibits a 20 kHz amplitude periodicity, with prominent lines attributed to the non-ideal extinction ratio of the EOIMs 39. (c) IGMs of the C-H-DCS and (d) its spectrum. The centerburst repetition rate is observed at 20 kHz, a 20th harmonic of its line spacing of 1 kHz, corresponding to the average mode amplitude in the spectrum increasing to $>$-60 dB. (e) IGMs of the H-H-DCS and (f) its spectrum. The centerburst repetition rate and the line spacing are both observed to be 1 kHz. The amplitude of the centerbursts is significantly increased to a value approaching 20 mVpp, corresponding to the spectral mode amplitude increased to close to -50 dB. (g) Evolution of time-domain SNR for the H-H-DCS and the C-DCS with the number of averages. On the log-log scale, the linear fitting models of the two datasets are given by fit 1 = 0.4997(4)x + 2.497(1) and fit 2 = 0.5075(4)x + 1.238(1). The slopes, approximating 0.5, indicate that the SNR is proportional to the square root of the number of averages, consistent with the principle of coherent averaging. The time-domain SNR is defined as the ratio of the centerburst amplitude to the standard deviation of a segment away from the centerburst 40.
[…]
Moreover, the two OFCs operated at repetition rates of 250 MHz and 249.98 MHz with a $\Delta f_{\mathrm{rep}}$ = 20 kHz, and carrier-envelope offset frequencies ($f_{\mathrm{ceo}}$) of 20 MHz and 33.5 MHz, respectively. To ensure long-term stability of the system, these $f_{\mathrm{ceo}}$ and $f_{\mathrm{rep}}$ were simply phase-locked to a rubidium atomic clock, with which the AWG was also timing synchronized (variable delay lines or RF phase shifters can be employed to align the optical pulses and modulation signals). Additionally, the high degree of mutual coherence essential for DCS was retrieved through the optical referencing and post-treatment techniques 19, 42, 43, which can be regarded as an original digital version of the adaptive scheme 44.
For the DCS experiments, an $\mathrm{H^{13}C^{14}N}$ gas cell (Wavelength References, HCN-13-10), with a nominal pressure of approximately 10 Torr and a path length of 16.5 cm, was selected as a representative sample. The two modulated combs were mixed and subsequently split into measurement and reference beams using a 50/50 coupler. The measurement beam traversed the $\mathrm{H^{13}C^{14}N}$ gas cell before detection by a biased detector (Thorlabs, DET01CFC), while the reference beam was directly detected. This reference channel is crucial for eliminating comb tooth fluctuations arising from the uneven gain caused by the error in $\phi_{TPM}$. IGMs from both detectors were low-pass filtered (6 MHz @3 dB) and simultaneously digitized by an oscilloscope at a sampling rate of 31.25 MS/s. The optical power incident on the detector in the measurement beam was $11.9~{}\mu W$ , $119~{}\mu W$, and $223~{}\mu W$ for the C-DCS, the C-H-DCS, and the H-H-DCS, respectively, limited by filter bandwidth and amplifier limitations (Supplementary Fig. S1).
IV Results and discussion
Fig. 6: Example $\mathrm{H^{13}C^{14}N}$ spectra relative to the center frequency of 193.5449 THz (1548.9555 nm) at a spectral resolution of 12.5 MHz, along with the corresponding weighted Voigt profile fits. (a) Recorded in 1.999 s (1999 averages) in the C-DCS. (b) Recorded in 1.999 s in the C-H-DCS. (c) Recorded in 5 ms (5 averages) in the H-H-DCS. (d) Recorded in 1.999 s in the H-H-DCS. The standard deviations of the four fitting residuals were calculated, with the residual noise in (d) is reduced by a factor of $\sim$20 compared to that in (a).
IV.1 Features of Interferograms
Figure 5(a–f) displays the IGMs generated in the three architectures and their corresponding spectra after post treatment. As anticipated, the centerburst repetition rate in the C-H-DCS is 20 times higher than that of the C-DCS, while maintaining the same spectral mode spacing and achieving an average enhancement of the mode amplitude by approximately 7 dB ($\mathrm{10lg\sqrt{20}}$).
As for the H-H-DCS, the centerburst amplitude is approximately 20 times greater than that in the C-DCS. This corresponds to an average enhancement of the mode amplitude by approximately 13 dB ($\mathrm{10lg20}$) in the spectrum. The noise floor in the spectrum also increases by about 2 dB, potentially due to higher relative intensity noise. Furthermore, in addition to the primary centerbursts, the IGMs of the H-H-DCS also show weak centerbursts with a 1/20 ms interval, which is attributed to imperfect transfer of the oscillation frequency due to errors in the driving signal levels. The mode superposition of these weak centerbursts caused the unflatness on the RF comb, i.e., an amplitude modulation with a period of 20 kHz, as shown in Figs. 5(e) and (f). Details regarding the higher-order components are presented in Supplementary Fig. S3.
A potential misconception might arise that the amplified centerburst amplitude necessitates a larger detector dynamic range. In fact, as mentioned in the principle section, this is a superposition resulting from the convolution of the additional pulse pairs due to low-pass filtering.
IV.2 Time-domain Signal-to-noise Radio
The evolution of the time-domain SNR for the H-H-DCS and the C-DCS with respect to the number of averages (or measurement time) is calculated in Fig. 5(g). The x-axis intercepts of the two linear fits exhibit a difference of 2.5575, indicating that this new architecture reduces the required number of averages or measurement time to achieve the same SNR by a factor of 361 (1/361 of the original). The y-axis intercepts show a difference of 1.259, indicating an approximate 18-fold enhancement in SNR for the same number of averages. These findings are consistent with the expected performance for a setup with m = 20.
IV.3 Transmission Spectrum
The enhancement in time-domain SNR is manifested in the transmission spectrum. Fig. 6 presents the acquired transmission spectra around the P7 absorption line of $\mathrm{H^{13}C^{14}N}$ using the three architectures, with a spectral point spacing of 12.5 MHz. Both architectures, particularly the H-H-DCS, demonstrate significantly reduced residual levels within the same measurement time compared to the C-DCS. Furthermore, the standard deviations of the fitting residual observed in a 5-ms measurement using H-H-DCS are comparable to those in a $\sim$ 2-s measurement using C-DCS.
IV.4 Discussion
The proof-of-concept results presented in this work can be further improved by combining with ultrahigh-repetition-rate comb sources, or when probing with finer spectral resolution. For example, when combined with 22-GHz microresonator combs 45 and operating at a spectral resolution of 500 MHz, compared to conventional DCS with the same resolution, an exceeding 1500-fold averaging-time reduction is expected. Furthermore, integrated lithium niobate electro-optic modulators operating at CMOS-compatible voltages may facilitate this implementation 46.
In wavelength regions where mature EOPMs are not readily available, this approach may can be achieved using nonlinear means. For example, under second-harmonic generation 47, a temporal pulse-phase profile of $[0,0,0,\pi/2]$ is anticipated to transform to the desired $[0,0,0,\pi]$, and a profile of $\pi\frac{s}{2m}k^{2}$ is anticipated to transform to the desired $\pi\frac{s}{m}k^{2}$.
For field-deployed DCS, the mitigation in coherence time inherently offers significant advantages. Further possible combination with ultra-low-noise optical frequency combs, such as the compact monolithic solid-state design 48 that provides a $\sim$ 100 Hz linewidth in a quasi-free-running state, complex and sophisticated locking systems may be greatly simplified. Moreover, the decoupling of the centerburst repetition rate and the $\Delta f_{\mathrm{rep}}$ (i.e. $\Delta f_{\mathrm{s}}$ ) also overcomes the limitation on the noise tracking bandwidth of post-treatment algorithms in free-running DCS, which was previously limited to half the IGM repetition rate (i.e. $\Delta f_{\mathrm{rep}}/2$) 43, 49, a restriction that can be roughly attributed to the fact that in conventional DCS, only a single centerburst within each IGM provides effective sampling.
Finally, it is worth emphasizing that the essence of this spectral mode enhancement lies in enabling the optical frequency comb or combs to output multiple special harmonic pulse trains with the same carrier-envelope offset frequency, but different carrier-envelope phases (non-linear slip), thereby enhancing the majority of the mode sets rather than enhancing a few and suppressing others. The scheme demonstrated in this paper, which employs TPM based on EOPMs outside the cavity, may not be the only method to achieve this SNR enhancement. For example, in some parametrically driven Kerr soliton microresonators, two soliton states with opposite phases owing to the $Z_{2}$-symmetry have been observed 50, and a soliton sequence with random phase bits of $[0,0,\pi,0]$ has been demonstrated 51.
V Conclusion
In summary, this work has opened up another effective way to enhance the SNR in DCS beyond coherent averaging, through coherent superposition of harmonic pulses. In our proof-of-concept experiments based on temporal phase modulation, we have successfully demonstrated an $\sim$ 18-fold SNR enhancement in a single acquisition, which is equivalent to an SNR gain achieved by coherent averaging over $>$300 periods in conventional DCS. This approach avoids the need for long-time averaging, thus enabling recovery of the inherent excellent temporal resolution of DCS and promising to relax stringent requirements on coherence time and stability of DCS systems. This could potentially have significant implications for future DCS research and applications.
Acknowledgments
This work was supported by National Key Research and Development Program of China under Grant2020YFC2200300. The authors thank Prof. Tang Li for providing the Menlo-250-WG-PM, Yatan Xiong for lending the EDFA, and Hao Li, Yinnan Chen, Zheng Liu, Yujia Ji, and Yanan Qu for their assistance with certain details.
Disclosures
The authors declare no conflicts of interest.
Data availability
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
Supplemental document
See Supplement 1 for supporting content.

### v4

@@
-Spectral Mode Enhancement in Coherent-harmonic Dual-comb Spectroscopy Enables Exceeding 300-fold averaging-time reduction
+Reducing Averaging Time in Dual-comb Spectroscopy via Phase-Patterned Higher-Repetition-Rate Pulses
 Abstract
-Dual-comb spectroscopy (DCS) is a novel Fourier-transform spectroscopy not relying on mechanical scanning and capable of simultaneously achieving high speed, high spectral resolution, and broad optical bandwidth.
-Nevertheless, it suffers from low signal-to-noise ratio (SNR) per single acquisition due to the dynamic range limitation of photodetectors imposed by the high-peak-power mode-locked pulses, making coherent averaging an essential means to improve SNR, at the price of compromising the exceptional time resolution and placing more stringent demands on mutual coherence and stability.
-In this study, a novel approach to enhance SNR by exploiting the spectral mode enhancement mechanism in coherent-harmonic pulses is demonstrated.
-As a proof-of-concept, two frequency combs with mode spacing of $\sim$12.5 MHz, operated at a 20th harmonic repetition rate of $\sim$250 MHz, are employed, demonstrating a $>$300-fold reduction in averaging time for comparable SNR in conventional DCS.
-This reduction is expected to be further enhanced through integration with ultra-high repetition rate combs like microresonator combs.
-This new approach promises both a recovery of the inherent high-speed capability and a mitigation of the coherence-time requirements, thereby making it possible to significantly facilitate subsequent DCS investigations and field deployments.
+Dual-comb spectroscopy (DCS) is a powerful Fourier-transform spectroscopic technique that provides high-speed, high-resolution, and broadband measurements without moving parts.
+However, the high peak power of mode-locked pulses limits the photodetector’s dynamic range, resulting in a low signal-to-noise ratio (SNR) per acquisition.
+While coherent averaging can improve SNR, it sacrifices temporal resolution and demands stringent system stability.
+Here, we introduce a novel concept to enhance SNR by using phase-patterned higher-repetition-rate combs.
+We reinterpret the self-imaging process of comb spectrum from a new perspective on mode interference among sub-pulse trains As a proof-of-concept, we densified two 250-MHz frequency combs to 12.5-MHz mode spacings via phase modulation and performed DCS on an $\mathrm{H^{13}C^{14}N}$ gas cell, and compared the results with an emulated conventional 12.5-MHz DCS, demonstrating a 17-fold increase in mode amplitude.
+This concept is expected to be combined with ultra-high repetition rate combs, such as microcombs, and thereby deployed in practical applications that typically require spectral sampling spacings from hundreds of MHz to GHz range.
 I Introduction
-Dual-comb spectroscopy (DCS) is a remarkable novel Fourier-transform spectroscopy technique powered by optical frequency comb (OFC) 1.
-Over the past decade, it has been increasingly recognized as an important spectroscopic tool in fields that require high speed, broad bandwidth, high frequency resolution, and high frequency accuracy, such as in investigations of physical turbulent environments 2, 3, 4, 5, 6 and chemical and biological dynamic processes 7, 8, 9, 10.
-By employing two coherent OFCs with a slight repetition rate difference ($\Delta f_{\mathrm{rep}}$), and inherently free from mechanical moving parts, DCS can capture an entire interferogram (IGM), yielding a spectrum with a full spectral resolution set by the repetition rate of the combs.
+Enabled by the unique properties of optical frequency combs (OFCs) 1, dual-comb spectroscopy (DCS) has emerged over the past decade as a powerful, next-generation Fourier-transform spectroscopy technique, offering an important tool for applications requiring high speed, broad bandwidth, high frequency resolution, and high frequency accuracy, such as in investigations of physical turbulent environments 2, 3, 4, 5, 6 and chemical and biological dynamic processes 7, 8, 9, 10.
+By employing two coherent OFCs with a slight repetition rate difference ($\Delta f_{\mathrm{rep}}$), and inherently free from mechanical moving parts, DCS can capture an interferogram (IGM), yielding a spectrum with a spectral resolution set by the repetition rate of the combs.
 The time frame for this capture is as short as $1/\Delta f_{\mathrm{rep}}$, and with large $\Delta f_{\mathrm{rep}}$ configurations, this leads to theoretical temporal resolution reaching microsecond or even nanosecond timescales 11, 12, 13.
 However, in typical practical DCS systems, the high peak power of mode-locked pulses often leads to photodetector saturation or nonlinear response, thus restricting the average power of the combs to a low level.
-Consequently, given the vast number of comb teeth inherent to broadband spectra, the power per comb tooth is typically limited to the nanowatt range 14, 15, ultimately resulting in extremely low SNR per single acquisition, which necessitates coherent averaging or multi-period acquisition to enhance the SNR 16, 15, 17, 12.
-Unfortunately, the SNR gain achieved through averaging scales proportionally to the square root of the number of averages.
+Consequently, given the vast number of comb teeth inherent to broadband spectra, the power per comb tooth is typically limited to the nanowatt range 14, ultimately resulting in extremely low SNR per single acquisition, When the comb source power is sufficient, this is referred to as the dynamic range limitation15, which necessitates coherent averaging or multi-period acquisition to enhance the SNR 16, 17, 12.
+Furthermore, the SNR gain achieved through averaging scales proportionally to the square root of the number of averages.
 For instance, to improve the SNR by a factor of 10, a hundred averages are needed.
@@
 However, the SNR enhancement offered by these two methods has been limited by experimental complexity or sampling bandwidth constraints, demonstrating averaging-time reduction of less than one order of magnitude.
-Another early approach is to stretch the pulses of the probe comb using a chirped fiber Bragg grating 19.
-This method can effectively reduce the dynamic range of an IGM; however, it has limited impact on the nonlinearity or saturation induced by pulses, as the detector’s response time significantly exceeds the pulse duration.
-More recently, the utilization of $\sim$ 1-GHz-repetition-rate combs 12, 20, 13, 21 offers higher average power and larger $\Delta f_{\mathrm{rep}}$.
-Nevertheless, a $\sim$ 1-GHz repetition rate had struck a delicate balance between measurement speed and spectral resolution 22.
-A higher repetition rate inherently implies a lower spectral resolution, which limits further increases in repetition rate.
-From this perspective, employing frequency combs analogous to certain harmonic mode-locked lasers 23, whose pulse repetition rate ($f_{\mathrm{rep}}$) operates on the harmonics of the laser cavity’s fundamental repetition rate, thus possibly featuring a high $f_{\mathrm{rep}}$ but a low mode spacing ($f_{\mathrm{ms}}$), may offer a promising avenue.
-On the one hand, this configuration can enable higher average power within the same dynamic range.
-On the other hand, the phenomenon that the pulse repetition rate exceeding the mode spacing likely points to an underlying mechanism of the harmonic pulses coherently enhancing their common spectral modes.
-If this mechanism can be effectively transferred to interferograms in DCS, i.e., generate multiple centerbursts within the time period of $1/\Delta f_{rep}$ to coherently enhance RF comb modes, it would be possible to achieve the desired SNR gain, thereby potentially refraining from extended IGMs coherent averaging.
-Here we find pulse temporal phase modulation (TPM) using electro-optical phase modulator (EOPM) provides an ideal means to realize such combs unaffected by supermode noise 32.
-This TPM has been employed for the comb densification technique 25, 26, 27, 28, 29, 30, 31 based on spectral self-imaging effect 24 over the past decade or so, which involves applying periodic TPM to the pulses, inducing evenly spaced ”self images” of the original comb lines on the spectrum, thereby densifying the comb.
-In 2016, N.B.
-Hébert et al.
-33 first introduced this densification technique in DCS and presented an enhanced sub-MHz spectral resolution from a 100-MHz comb, using a pseudo-random binary sequence (PRBS) modulation.
-In 2024, C.
-Q.-Galan et al.
-34 extended this method to 1-GHz gain-switched OFCs, achieving $\sim$ 3.9 MHz resolution.
-Despite this, both of these two studies primarily focused on the spectral resolution enhancement, merely treating such combs as conventional OFCs with small mode spacing.
-The underlying harmonic-enhancement mechanism and the descriptions of such special dual-comb have not been elucidated.
-In this work, we derived a coherent-harmonic-pulse superposition model in the frequency domain to elucidate this mechanism, and presented the time-domain and equivalent frequency-domain descriptions of two architectures for this special DCS, with a focus on their respective SNR gains.
-To align with its intuitive physical picture and distinguish it from conventional dual-comb spectroscopy (C-DCS), we term this approach coherent-harmonic-enhanced dual-comb spectroscopy (CH-EDCS).
-As a proof of concept, we perform transmission spectrum measurements of a hydrogen cyanide ($\mathrm{H^{13}C^{14}N}$) gas cell the two proposed architectures and, for comparison, a C-DCS architecture.
-The experiments were based on two commercial 250 MHz combs, which were configured using TPM and a pulse-picking technique 41 to operate at a spectral resolution of 12.5 MHz.
-As a result, we achieved an SNR gain of up to $\sim$ 18 times per single acquisition, equivalent to a $>$300-fold reduction in averaging time, demonstrating the significant potential for SNR enhancement of this approach.
+More recently, the utilization of $\sim$ 1-GHz-repetition-rate combs 12, 19, 13, 20 offers higher average power and larger $\Delta f_{\mathrm{rep}}$.
+Nevertheless, a $\sim$ 1-GHz repetition rate had struck a delicate balance between measurement speed and spectral bin spacing 21.
+A higher repetition rate inherently implies a lower spectral bin spacing, which limits further increases in repetition rate.
+Over the past two decades, a comb densification technique based on the spectral self-imaging effect (also known as the spectral Talbot effect) 22 has been developed 23, 24, 25, 26, 27, 28.
+This technique applies periodic temporal-phase modulation (TPM) to the comb pulses, thereby inducing evenly spaced ”self images” of the original comb modes and leading to the densification of the comb 29, 30, 31.
+In this paper, we model the modulated pulse train as a collection of multiple sub-pulse trains and reinterpret the self-imaging process of comb spectrum from a new perspective on mode interference among these trains.
+This perspective is analogous to the supermode analysis in harmonic mode-locking32, 33.
+Subsequently, we intuitively demonstrate how multiple sub-pulse trains synergistically enhance their common modes.
+This shows that mode energy can be increased using higher-repetition-rate pulses within the same detector dynamic range without reducing spectral bin spacing, thereby providing an effective solution to the SNR limitation in DCS.
+Building on this framework, we provide a complete description for two architectures for this special form of DCS.
+As a proof-of-concept, we densify two close-to-250 MHz frequency combs via periodic TPM and measured the transmission spectrum of a hydrogen cyanide ($\mathrm{H^{13}C^{14}N}$) gas cell at a spectral sampling bin spacing of 12.5 MHz.
+Compared to an emulated conventional 12.5-MHz DCS, we demonstrate a 17-fold increase in mode amplitude and an SNR gain of up to $\sim$16.7 times per single acquisition in the non-shot-noise-limited regime, the latter corresponding to a 290-fold reduction in averaging time.
+This average time reduction is expected to reduce to 17-fold in the shot-noise-limited regime.
+To distinguish it from conventional dual-comb spectroscopy (C-DCS), we term this approach phase-patterned higher-repetition-rate-pulse enhanced dual-comb spectroscopy (PPHE-DCS).
+II Principles
+In this section, we first present the spectral behavior of phase-patterned higher-repetition-rate pulses from a perspective of mode interference among sub-pulse trains.
+Then, two DCS architectures for this PPHE-DCS are studied: one employing a conventional optical frequency comb (C-OFC) and a phase-patterned higher-repetition-rate optical frequency comb (PH-OFC), termed C-PH-DCS; the other employing two PH-OFCs, termed PH-PH-DCS.
+For clarity, we use a pulse-multiplexing number of m = 4 and a TPM sequence of $\pi,0,0,0]$ as illustrative examples.
+II.1 Spectrum of Phase-Patterned Higher-Repetition-Rate Pulses
+As a example, when phase modulation an OFC with a repetition frequency of $4f_{\mathrm{ms}}$ with a $f_{\mathrm{ms}}$-periodic sequence of $\pi,0,0,0]$, the original pulse train can be decomposed into a collection of four sub-pulse trains, each with a repetition frequency of $f_{\mathrm{ms}}$, as illustrated in Fig.
+1(a).
+The pulses within each sub-pulse train are still fully coherent, thereby each sub-pulse train corresponding to a sub-spectrum with mode spacing of $f_{\mathrm{ms}}$ (Fig.
+1(b)).
+According to the linear property of the Fourier transform, the spectrum of the PH-OFC or the interferogram at frequency $\omega_{j}$ is a vector superposition of the 4 sub-modes at that frequency, given by $E_{j}=\sum_{k=1}^{4}a_{0}\exp{(\omega_{j}t+\phi_{k,j})}$, where $a_{0}$ is the amplitude of the sub-modes.
+Considering the influence of the relative time delay among these sub-pulse trains and the TPM, the phase spectrum of the kth-sub-spectrum can be expressed as
+$\phi_{k,j}=-\omega_{j}t_{k}+\phi_{TPM,k}=-\frac{j}{4}(k-1)\cdot 2\pi+\phi_{TPM,k}.$ (1)
+The first term in the formula, $\omega_{j}t_{k}$, arises from the contribution of the time shift, where $\omega_{j}=j\cdot 2\pi f_{\mathrm{ms}}$ is the angular frequency of the $j_{th}$ mode, $t_{k}=(k-1)/4f_{\mathrm{ms}}$ is the time delay of kth-sub-pulse train relative to the first sub-pulse train; the second term, $\phi_{TPM,k}$, is the phase shift added to the kth-train through the TPM.
+The superposition process and results are illustrated in Figs.
+1(b) and (c), with vector sum $E_{j}=2a_{0}\exp{(\omega_{j}t+\varphi_{j})}$ where $\varphi_{j}$ is a periodic sequence $\pi,\pi,\pi,0]$.
+This indicates that the spectrum of the phase-patterned higher-repetition-rate pulses train has the same mode spacing but 2-fold the amplitude of the sub-spectrum (which directly corresponds to a 4-fold increase in pulses, i.e., a 4-fold increase in optical power), and exhibits a periodic multilevel phase spectrum of $\pi,\pi,\pi,0]$.
+This form of spectrum can be viewed as the result of interleaving four independent modesets, within which the mode spacing is $f_{\mathrm{ms}}$ and the modes are perfectly phase-coherent.
+In practice, a deviation in the phase modulation level leads to inconsistent interference of modes at different frequencies when all the sub-spectra are superimposed (imaging that the rotation angle of the blue arrow in Fig.
+1(c) is not exactly 180 degrees), which in turn causes periodic mode-to-mode-amplitude variations.
+This also corresponds to the more general case of using an arbitary signal and the resulting spectrum can still be viewed as the result of interleaving four independent modesets.
+When the period of the modulation signal is greater than 4, to obtain a consistent constructive interference, one can use either the quadratic function form indicated by the so-called ”Talbot condition” 34 or a binary pseudo- random signal26.
+The equation of the former is expressed as
+$\phi_{n;s,m}=\pi\frac{s}{m}k^{2}(mod~2\pi),$ (2)
+where s, m are mutually prime integer numbers, and k is the pulse number.
+Correspondingly, the periodic multilevel phase profile of the total spectrum is $\varphi_{k;p,q}=-\pi\frac{p}{q}j^{2}(mod~2\pi),$ where $q=m$, j is the mode number, and $p\neq s$ in general.
 Fig.
-1: Concepts of conventional DCS (C-DCS) and C-H-DCS.
+1: Illustration of the process of the spectral mode superposition in phase-patterned sub-pulse trains.
+(a) Time domain description.
+(b) The equivalent frequency-domain description.
+The j represents the spectral mode index (derived from $\omega_{j}=j\cdot 2\pi f_{\mathrm{ms}}$).
+The $\pi$ phase of the first spectrum originates from the $\pi$ phase shift in the time domain; the phase slopes of the 2nd to 4th spectra are caused by the phase shift of $\omega_{j}t_{k}$ arising from the temporal shift of pulses.
+(c) Illustration of vector superposition of the modes of the four sub-pulse trains.
+The colored linear arrows represent the mode from the kth pulse train at $\omega_{j}$, while the blue rotating arrows indicate the temporal $\pi$-phase shift on the k = 1 pusle train, which rotates all its spectral modes by $\pi$.
+The black arrows represent the total mode vectors after superposition, which have twice the amplitude of the sub-modes and exhibit a periodic phase profile of $\pi,-\pi,-\pi,0]$.
+In the absence of the $\pi$-shift, only the modes of $j=4n$ $(n\in\mathbb{N})$ will be present, while the remaining modes will undergo destructive interference.
+Consequently, the mode spacing is consistent with the temporal pulse repetition rate, at $4f_{\mathrm{ms}}$, which aligns with classic OFCs.
+Fig.
+2: Concepts of conventional DCS (C-DCS) and C-PH-DCS.
 (a) Time-domain description of C-DCS.
 The pulse train (red) of comb 1 progressively steps across another pulse train (blue) with a slight repetition rate difference $\Delta f_{\mathrm{s}}$, generating an interferogram with a single centerburst in a period of $1/\Delta f_{\mathrm{s}}$.
-When the temporal interval between the two pulses is large, effective interference does not occur.(b) The equivalent frequency-domain description of the C-DCS.
+When the temporal interval between the two pulses is large, effective interference does not occur.
+(b) The equivalent frequency-domain description of the C-DCS.
 The beating of two combs produces an RF comb in the radio-frequency domain.
-(c) Time-domain description of the C-H-DCS.
-The pulse of the conventional optical frequency comb (C-OFC) with a pulse repetition rate of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$, steps across the four pulses of the coherent-harmonic optical frequency comb (H-OFC), which have phases of $[\pi,0,0,0]$ and a pulse repetition rate of $4f_{\mathrm{ms}}$.
-This generates four centerbursts within an interferogram period of $1/\Delta f_{\mathrm{s}}$, which inherit the $[\pi,0,0,0]$ phases.
-(d) The equivalent frequency-domain description of the C-H-DCS.
+(c) Time-domain description of the C-PH-DCS.
+The pulse of the conventional optical frequency comb (C-OFC) with a pulse repetition rate of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$, steps across the four pulses of the phase-patterned higher-repetition-rate optical frequency comb (PH-OFC), which have phases of $\pi,0,0,0]$ and a pulse repetition rate of $4f_{\mathrm{ms}}$.
+This generates four centerbursts within an interferogram period of $1/\Delta f_{\mathrm{s}}$, which inherit the $\pi,0,0,0]$ phases.
+(d) The equivalent frequency-domain description of the C-PH-DCS.
 Two combs with mode spacing of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$ and $f_{\mathrm{ms}}$ respectively beat together, producing an RF comb in the radio-frequency domain.
-This RF comb inherits the $[0,-\pi,-\pi,-\pi]$ periodic multilevel phase profile of the H-OFC and exhibits twice the line amplitude compared to the C-DCS.
-II Principles
-In this section, two DCS architectures for this CH-EDCS concept are studied: one termed C-H-OFC, employing a coherent-harmonic optical frequency comb (H-OFC) realized by TPM and a conventional optical frequency comb (C-OFC); the other termed H-H-DCS, employing two H-OFCs.
-For clarity, we use a harmonic number of m = 4 and a TPM sequence of $[\pi,0,0,0]$ as illustrative examples.
-II.1 C-H-DCS
+This RF comb inherits the $\pi,\pi,\pi,0]$ periodic multilevel phase profile of the PH-OFC and exhibits twice the line amplitude compared to the C-DCS.
+Fig.
+3: Concept of PH-PH-DCS.
+(a) Time-domain description of the PH-PH-DCS.
+The PH-OFC 1 and PH-OFC 2 emit two pulse trains with repetition rate of $4(f_{\mathrm{ms}}-\Delta f_{\mathrm{s}})$ and $4f_{\mathrm{ms}}$ respectively, both with pulse phases of $\pi,0,0,0]$.
+Interference of the two pulse trains within a time period of $1/\Delta f_{\mathrm{s}}$ produces four distinct centerbursts: one at the fundamental oscillation frequency (purple), exhibiting a 4-fold amplitude enhancement over conventional dual-comb spectroscopy (C-DCS); and three additional centerbursts at higher-order frequencies (pink), which are suppressed by low-pass filtering in practical implementations.
+(b) The equivalent frequency-domain description of the PH-PH-DCS.
+Two PH-OFCs with mode spacing of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$ and $f_{\mathrm{ms}}$ respectively beat together, producing an RF comb in the radio-frequency domain which has a uniform phase profile and exhibits four times the line amplitude compared to the C-DCS.
+Note that, the uniform phase profile requires the phase profiles of the two PH-OFCs to be accurately aligned.
+Fig.
+4: Experimental setup.
+Rb clock: rubidium clock; OFC: optical frequency comb; AWG: arbitrary waveform generator; RFA: RF amplifiers; EOPM/EOIM: electro-optical phase/intensity modulator; OC: optical coupler; PD: photodetector; LPF: low-pass filter; BPF: band-pass filter; EDFA: erbium-doped fiber amplifier; CW: continuous wave laser; DWDM: dense wavelength-division multiplexer.
+II.2 C-PH-DCS
 As illustrated in Figs.
-1 (c) and (d), C-H-DCS utilizes a H-OFC whose repetition rate is approximately equal to 4 times of the repetition rate of the C-OFC (or, more generally, any integer multiple m).
-The pulses of comb 1, after stepping across the first pulse of the H-OFC, subsequently scan its 2nd to 4th pulses, thus generating four centerbursts instead of a single one on the interferogram within the same time period.
+2(c) and (d), C-PH-DCS utilizes a PH-OFC whose repetition rate is approximately equal to 4 times of the repetition rate of the C-OFC (or, more generally, any integer multiple m).
+The pulses of comb 1, after stepping across the first pulse of the PH-OFC, subsequently scan its 2nd to 4th pulses.
+Consequently, four centerbursts are generated within a single period of $1/\Delta f_{\mathrm{s}}$, as opposed to the single one generated in C-DCS (Fig.
+2(a)).
 Additionally, according to the principle of linear optical sampling 35, 36, the expression for the interference intensity can be written as
@@
 For each centerburst, the $\phi_{TPM}$ is constant.
-Consequently, the TPM imposed on the pulses of H-OFC is finally mapped onto the four centerbursts.
-Fig.
-2: Illustration of the process of the spectral mode enhancement by the harmonic pulse trains in the frequency domain.
-The j represents the spectral mode index (derived from $\omega_{j}=j\cdot 2\pi f_{\mathrm{ms}}$).
-The colored linear arrows with indexes of k = 1, 2, 3, 4 represent the mode vectors from the kth-sub-harmonic pulse train at the frequency of $\omega_{j}$, while the blue rotating arrows indicate the temporal phase modulation (TPM), which rotates all the modes of the 1st-sub-harmonic pulse train (k = 1) by $\pi$.
-In the absence of this TPM, only the modes of $j=4n$ $(n\in\mathbb{Z}^{+})$ will be present, while the remaining modes will undergo destructive interference, consistent with conventional expectation.
-The black arrows represent the total mode vectors after superposition, which have twice the amplitude of the sub-modes and exhibit a periodic phase profile of $[-\pi,-\pi,-\pi,0]$.
-Fig.
-3: Concept of H-H-DCS.
-(a) Time-domain description of the H-H-DCS.
-The H-OFC 1 and H-OFC 2 emit two pulse trains with repetition rate of $4(f_{\mathrm{ms}}-\Delta f_{\mathrm{s}})$ and $4f_{\mathrm{ms}}$ respectively, both with pulse phases of $[\pi,0,0,0]$.
-Interference of the two pulse trains within a time period of $1/\Delta f_{\mathrm{s}}$ produces four distinct centerbursts: one at the fundamental oscillation frequency (purple), exhibiting a 4-fold amplitude enhancement over conventional dual-comb spectroscopy (C-DCS); and three additional centerbursts at higher-order frequencies (pink), which are suppressed by low-pass filtering in practical implementations.
-(b) The equivalent frequency-domain description of the H-H-DCS.
-Two H-OFCs with mode spacing of $f_{\mathrm{ms}}-\Delta f_{\mathrm{s}}$ and $f_{\mathrm{ms}}$ respectively beat together, producing an RF comb in the radio-frequency domain which has a uniform phase profile and exhibits four times the line amplitude compared to the C-DCS.
-Note that, the uniform phase profile requires the phase profiles of the two H-OFCs to be accurately aligned.
-The spectral behavior of the H-OFC and the interferogram with 4 centerbursts can be elucidated through the interference of the harmonic pulses.
-The original pulse train is decomposed into a collection of four sub-harmonic pulse trains, each with a period of $1/f_{\mathrm{ms}}$ ($1/\Delta f_{\mathrm{s}}$ for the interferogram) and corresponding to a sub-spectrum with mode spacing of $f_{\mathrm{ms}}$.
-According to the linear property of the Fourier transform, the spectrum of the H-OFC or the interferogram at frequency $\omega_{j}$ is a vector superposition of the 4 sub-modes at that frequency, given by $E_{j}=\sum_{k=1}^{4}a_{0}\exp{(\omega_{j}t+\phi_{k,j})}$, where $a_{0}$ is the amplitude of the sub-modes.
-Considering the influence of the relative time delay among these harmonic trains and the TPM, the phase spectrum of the kth-sub-spectrum can be expressed as
-$\phi_{k,j}=-\omega_{j}t_{k}+\phi_{TPM,k}=-\frac{j}{4}(k-1)\cdot 2\pi+\phi_{TPM,k}.$ (2)
-The first term in the formula, $-\omega_{j}t_{k}$, arises from the contribution of the Fourier time shift, where $\omega_{j}=j\cdot 2\pi f_{\mathrm{ms}}$ is the angular frequency of the $j_{th}$ mode, $t_{k}=(k-1)/4f_{\mathrm{ms}}$ is the time delay of kth-sub-harmonic pulse train relative to the first sub-harmonic pulse train; the second term, $\phi_{TPM,k}$, is the phase shift added to the kth-train through the TPM.
-The superposition process and results are illustrated in Fig.
-2, with vector sum $E_{j}=2a_{0}\exp{(\omega_{j}t+\varphi_{j})}$ where $\varphi_{j}$ is a periodic sequence $[-\pi,-\pi,-\pi,0]$.
-This indicates that the spectrum of the original pulse train has the same mode spacing but twice the amplitude of the sub-spectrum, and exhibits a periodic multilevel phase spectrum of $[-\pi,-\pi,-\pi,0]$.
-Thus, an SNR gain of 2 can be achieved without coherent averaging 4 IGMs (spanning a total time of $4/\Delta f_{\mathrm{s}}$ in the C-DCS).
-For a more general case, the number 4 in the aforementioned equations corresponds to m, and the corresponding SNR gain is $\sqrt{m}$.
-Moreover, it is worth noting that if a flat gain is required, an equation in the quadratic function form indicated by the so-called ”Talbot condition” 37 can be used, which is expressed as
-$\phi_{n;s,m}=\pi\frac{s}{m}k^{2}(mod~{}2\pi),$ (3)
-where s, m are mutually prime integer numbers, and k is the pulse number.
-Correspondingly, the periodic multilevel phase profile of the total spectrum is $\varphi_{k;p,q}=-\pi\frac{p}{q}j^{2}(mod~{}2\pi),$ where $q=m$, j is the mode number, and $p\neq s$ in general.
-Whereas, if a reference channel for spectral normalization in DCS configuration is available, the majority of periodic signal can be set.
-II.2 H-H-DCS
+Consequently, the TPM imposed on the pulses of PH-OFC is finally mapped onto the four centerbursts.
+The spectral behavior of the interferogram with four phase-patterned centerbursts can be also deduced from the spectral mode superposition process as illustrated in Fig.
+1.
+Consequently, the C-PH-DCS (Fig.
+2(d)) provides a 2-fold amplitude-enhancement in spectral mode compared to the C-DCS (Fig.
+2(b)), all while preserving the spectral sampling spacing.
+For a more general case, the number 2 aforementioned corresponds to $\sqrt{m}$, where m is defined as the multiplexing factor of the pulses.
+II.3 PH-PH-DCS
 As illustrated in Fig.
-3, C-H-DCS utilizes two H-OFCs with pulse repetition rates of $4f_{\mathrm{ms}}$ and $4(f_{\mathrm{ms}}-\Delta f_{\mathrm{s}})$.
+3, PH-PH-DCS utilizes two PH-OFCs with pulse repetition rates of $4f_{\mathrm{ms}}$ and $4(f_{\mathrm{ms}}-\Delta f_{\mathrm{s}})$.
 In the time domain, similar to a C-DCS with the same repetition rate difference of $4\Delta f_{\mathrm{s}}$, this architecture generates 4 centerbursts within a time period of $1/\Delta f_{\mathrm{s}}$.
-Due to this repetition rate difference, the phase-inverted pulses in the two pulse trains gradually slip relative to each other, resulting in distinct in-pulse-pair phase configurations: $[\pi,0,0,0]-[\pi,0,0,0]$, $[0,\pi,0,0]-[\pi,0,0,0]$, $[0,0,\pi,0]-[\pi,0,0,0]$, and $[0,0,0,\pi]-[\pi,0,0,0]$, respectively, in the time domain regions corresponding to the four centerbursts.
+Due to this repetition rate difference, the phase-inverted pulses in the two pulse trains gradually slip relative to each other, resulting in distinct in-pulse-pair phase configurations: $\pi,0,0,0]-[\pi,0,0,0]$, $\pi,0,0]-[\pi,0,0,0], [0,0,$$0,\pi,0]-[\pi,0,0,0$], and $0,0,0,\pi]-[\pi,0,0,0]$, respectively, in the time domain regions corresponding to the four centerbursts.
 Consequently, the $\phi_{TPM}$ in Eq.
-1, is constantly 0 for the first centerburst; while for the other three centerbursts, the values are periodic $[\pi,\pi,0,0],[\pi,0,\pi,0]$, and $[\pi,0,0,\pi]$, respectively.
+3, is constantly 0 for the first centerburst; while for the other three centerbursts, the values are periodic $\pi,\pi,0,0],[\pi,0,\pi,0]$, and $\pi,0,0,\pi]$, respectively.
 These time-varying phases (related to the pulse pair index, i.e., time) increase the oscillation frequency of the interference electric field, shifting the 2nd to 4th centerbursts to higher-order components, which can be filtered out by an electronic low-pass filter in practical implementation.
@@
 Distinctly, the fundamental centerburst originates from the contribution of 4 times the number of pulse pairs, whose interference intensity is given by $I^{\prime}(n)\sim\sum_{n=1}^{4n_{0}}\langle|E_{1}(t-n\tau/4)||E_{2}(t)|\cos\Delta\omega_{c}t\rangle$, where $n_{0}$ is the number of pulse pairs contributing to the centerburst in the C-DCS illustrated in Fig.
-1(a).
+2(a).
 Consequently, following convolution with the detector response and low-pass filtering, the amplitude of this centerburst is also enhanced by a factor of 4.
-In the frequency domain, two H-OFCs with 2-fold mode amplitude enhancement and periodic multilevel phase spectra of $[-\pi,-\pi,-\pi,0]$, beat together to yield an RF comb.
-Since the amplitude of a heterodyne beat note is the product of the amplitudes of the two beating components, and the phase is the phase difference between them, the modes of this RF comb therefore exhibit a 4-fold enhancement in amplitude and possess a uniform phase, which precisely corresponds to the single fundamental centerburst with a 4-fold increase in amplitude in the time domain.
-Thus, this architecture leads to a 4-fold SNR gain.
-For a more general case, this number 4 corresponds to m, and is equivalent to an SNR gain brought by coherent averaging spanning a total time of $m^{2}$-fold.
-Interestingly, the aforementioned IGM characteristics can be changed (the SNR gain is not changed) when the two TPM signals use the quadratic form described in Eq.
+In the frequency domain, two PH-OFCs with 2-fold mode amplitude enhancement compared to the C-OFC and periodic multilevel phase spectra of $\pi,\pi,\pi,0]$, beat together to yield an RF comb.
+Since the amplitude of a heterodyne beat note is the product of the amplitudes of the two beating components, and the phase is the phase difference between them, the modes of this RF comb therefore exhibit a 4-fold enhancement in amplitude and possess a uniform phase, which precisely corresponds to the single fundamental centerburst with a 4-fold enhancement in amplitude in the time domain.
+For a more general case, the amplitude enhancement is m-fold.
+A potential misconception might arise that the amplified centerburst amplitude necessitates a larger detector dynamic range.
+In fact, this is a superposition resulting from the convolution of the additional pulse pairs due to low-pass filtering.
+Interestingly, the aforementioned IGM characteristics can be changed when the two TPM signals use the quadratic form described in Eq.
 2 with the same m but different s values.
-The generated RF comb can exhibit quadratic periodic multilevel phase profiles of $-\pi\frac{p_{1}-p_{2}}{m}j^{2}$.
-According to the temporal self-imaging effect 38, 30, the fundamental centerburst on the IGM will then be redistributed, similar to that in C-H-DCS.
+The generated RF comb can exhibit quadratic periodic multilevel phase profiles of $\pi\frac{p_{1}-p_{2}}{m}j^{2}$.
+According to the temporal self-imaging effect 37, 27, the fundamental centerburst on the IGM will then be redistributed to m/2 copies, similar to that in C-PH-DCS.
+A similar effect is expected when using two PRBS signals with different patterns.
+This will help to alleviate the dynamic range pressure on components within the detection chain other than the detector.
 III Experiment
+[…]
+The two OFCs operated at repetition rates of 250 MHz and 249.98 MHz with a $\Delta f_{\mathrm{rep}}$ = 20 kHz, and carrier-envelope offset frequencies ($f_{\mathrm{ceo}}$) of 20 MHz and 33.5 MHz, respectively.
+To ensure long-term stability of the system, these $f_{\mathrm{ceo}}$ and $f_{\mathrm{rep}}$ were simply phase-locked to a rubidium atomic clock, with which the AWG was also timing synchronized.
+Variable delay lines or RF phase shifters can be employed to align the optical pulses and modulation signals.
+The high degree of mutual coherence essential for DCS was retrieved through the optical referencing and post-treatment techniques40.
+A continuous wave (CW) laser is used as an intermediate oscillator 41 and the self-correction algorithm based on cross-correlation function 42, 43 are implemented.
+Furthermore, operating at spectral bin spacing of 12.5 MHz leads to a limited available spectral bandwidth.
+According to the DCS non-aliasing condition 14, the available bandwidth satisfies $\Delta\nu\leq f_{\mathrm{ms}}^{2}/{2\Delta f_{\mathrm{s}}}$, which yields 78.125 GHz (0.626 nm at 1550 nm) here for $f_{\mathrm{ms}}$ = 12.5 MHz and $\Delta f_{\mathrm{s}}$ = 1 kHz.
+Therefore, optical band-pass filters (BPFs) were employed to narrow the spectrum, specifically two 1.3-nm @ 25 dB filters (BPF 1) and a 0.3-nm @ 12 dB filter (BPF 2).
+Because most of the optical power was lost during filtering, two EDFAs were used to partially compensate.
+The optical power incident on the detector in the measurement beam was $1.9~\mu W, 11$$19~\mu W$, and $223~\mu W$ for the C-DCS, the C-PH-DCS, and the PH-PH-DCS.
+For the DCS measurement, an $\mathrm{H^{13}C^{14}N}$ gas cell (Wavelength References, HCN-13-10) with a nominal pressure of approximately 10 Torr and a path length of 16.5 cm was selected as a representative sample.
+In addition to the measurement beam, a reference beam was directly detected for baseline normalization.
+Both beams were detected by two InGaAs biased photodetectors (Thorlabs DET01CFC, 1 GHz bandwidth, no internal amplification).
+The photocurrent was low-pass filtered (6 MHz @ 3 dB) and then converted to a voltage signal by terminating the detector directly into the 50 $\Omega$ input of a Tektronix MSO54 oscilloscope.
+The signal was then digitized at a sampling rate of 31.25 MS/s with 12-bit resolution.
+IV Results and discussion
 Fig.
-4: Sketch of experimental setup.
-Rb clock: rubidium clock; OFC: optical frequency comb; AWG: arbitrary waveform generator; RFA: RF amplifiers; EOPM/EOIM: electro-optical phase/intensity modulator; OC: optical coupler; PD: photodetector; LPF: low-pass filter.
-Fig.
-5: Comparison of interferograms (IGMs) and signal-to-noise ratio (SNR) for the three architectures.
+5: Comparison of interferograms (IGMs) and spectrum for the three architectures.
 (a) IGMs of the C-DCS and (b) its spectrum (partial).
 The centerburst repetition rate is 1 kHz, matching the spectral line spacing, as expected for conventional DCS.
-The spectrum exhibits a 20 kHz amplitude periodicity, with prominent lines attributed to the non-ideal extinction ratio of the EOIMs 39.
-(c) IGMs of the C-H-DCS and (d) its spectrum.
-The centerburst repetition rate is observed at 20 kHz, a 20th harmonic of its line spacing of 1 kHz, corresponding to the average mode amplitude in the spectrum increasing to $>$-60 dB.
-(e) IGMs of the H-H-DCS and (f) its spectrum.
+The spectrum exhibits a 20 kHz amplitude periodicity, with prominent lines attributed to the non-ideal extinction ratio of the EOIMs 38.
+(c) IGMs of the C-PH-DCS and (d) its spectrum.
+The centerburst repetition rate is observed at 20 kHz, a 20th harmonic of its line spacing of 1 kHz, corresponding to the average mode amplitude in the spectrum increasing to >>-60 dB.
+(e) IGMs of the PH-PH-DCS and (f) its spectrum.
 The centerburst repetition rate and the line spacing are both observed to be 1 kHz.
 The amplitude of the centerbursts is significantly increased to a value approaching 20 mVpp, corresponding to the spectral mode amplitude increased to close to -50 dB.
-(g) Evolution of time-domain SNR for the H-H-DCS and the C-DCS with the number of averages.
-On the log-log scale, the linear fitting models of the two datasets are given by fit 1 = 0.4997(4)x + 2.497(1) and fit 2 = 0.5075(4)x + 1.238(1).
-The slopes, approximating 0.5, indicate that the SNR is proportional to the square root of the number of averages, consistent with the principle of coherent averaging.
-The time-domain SNR is defined as the ratio of the centerburst amplitude to the standard deviation of a segment away from the centerburst 40.
-[…]
-Moreover, the two OFCs operated at repetition rates of 250 MHz and 249.98 MHz with a $\Delta f_{\mathrm{rep}}$ = 20 kHz, and carrier-envelope offset frequencies ($f_{\mathrm{ceo}}$) of 20 MHz and 33.5 MHz, respectively.
-To ensure long-term stability of the system, these $f_{\mathrm{ceo}}$ and $f_{\mathrm{rep}}$ were simply phase-locked to a rubidium atomic clock, with which the AWG was also timing synchronized (variable delay lines or RF phase shifters can be employed to align the optical pulses and modulation signals).
-Additionally, the high degree of mutual coherence essential for DCS was retrieved through the optical referencing and post-treatment techniques 19, 42, 43, which can be regarded as an original digital version of the adaptive scheme 44.
-For the DCS experiments, an $\mathrm{H^{13}C^{14}N}$ gas cell (Wavelength References, HCN-13-10), with a nominal pressure of approximately 10 Torr and a path length of 16.5 cm, was selected as a representative sample.
-The two modulated combs were mixed and subsequently split into measurement and reference beams using a 50/50 coupler.
-The measurement beam traversed the $\mathrm{H^{13}C^{14}N}$ gas cell before detection by a biased detector (Thorlabs, DET01CFC), while the reference beam was directly detected.
-This reference channel is crucial for eliminating comb tooth fluctuations arising from the uneven gain caused by the error in $\phi_{TPM}$.
-IGMs from both detectors were low-pass filtered (6 MHz @3 dB) and simultaneously digitized by an oscilloscope at a sampling rate of 31.25 MS/s.
-The optical power incident on the detector in the measurement beam was $11.9~{}\mu W$ , $119~{}\mu W$, and $223~{}\mu W$ for the C-DCS, the C-H-DCS, and the H-H-DCS, respectively, limited by filter bandwidth and amplifier limitations (Supplementary Fig.
-S1).
-IV Results and discussion
-Fig.
-6: Example $\mathrm{H^{13}C^{14}N}$ spectra relative to the center frequency of 193.5449 THz (1548.9555 nm) at a spectral resolution of 12.5 MHz, along with the corresponding weighted Voigt profile fits.
-(a) Recorded in 1.999 s (1999 averages) in the C-DCS.
-(b) Recorded in 1.999 s in the C-H-DCS.
-(c) Recorded in 5 ms (5 averages) in the H-H-DCS.
-(d) Recorded in 1.999 s in the H-H-DCS.
-The standard deviations of the four fitting residuals were calculated, with the residual noise in (d) is reduced by a factor of $\sim$20 compared to that in (a).
 IV.1 Features of Interferograms
-Figure 5(a–f) displays the IGMs generated in the three architectures and their corresponding spectra after post treatment.
-As anticipated, the centerburst repetition rate in the C-H-DCS is 20 times higher than that of the C-DCS, while maintaining the same spectral mode spacing and achieving an average enhancement of the mode amplitude by approximately 7 dB ($\mathrm{10lg\sqrt{20}}$).
-As for the H-H-DCS, the centerburst amplitude is approximately 20 times greater than that in the C-DCS.
+Figure 5(a–f) displays the raw IGMs generated in the three architectures and their corresponding spectra after post treatment.
+As anticipated, the centerburst repetition rate in the C-PH-DCS is 20 times higher than that of the C-DCS, while maintaining the same spectral mode spacing and achieving an average enhancement of the mode amplitude by approximately 7 dB ($\mathrm{10lg\sqrt{20}}$).
+As for the PH-PH-DCS, the centerburst amplitude is approximately 20 times greater than that in the C-DCS.
 This corresponds to an average enhancement of the mode amplitude by approximately 13 dB ($\mathrm{10lg20}$) in the spectrum.
-The noise floor in the spectrum also increases by about 2 dB, potentially due to higher relative intensity noise.
-Furthermore, in addition to the primary centerbursts, the IGMs of the H-H-DCS also show weak centerbursts with a 1/20 ms interval, which is attributed to imperfect transfer of the oscillation frequency due to errors in the driving signal levels.
+The noise floor in the spectrum also increases by about 2 dB.
+Furthermore, in addition to the primary centerbursts, the IGMs of the PH-PH-DCS also show weak centerbursts with a 1/20 ms interval.
+This is attributed to an error in the modulation signal level, which prevented the original centerbursts at these adjacent locations from being completely transferred to high-frequency components (Fig.
+3(a)), leaving a small residue at the fundamental oscillation frequency.
 The mode superposition of these weak centerbursts caused the unflatness on the RF comb, i.e., an amplitude modulation with a period of 20 kHz, as shown in Figs.
 5(e) and (f).
-Details regarding the higher-order components are presented in Supplementary Fig.
-S3.
-A potential misconception might arise that the amplified centerburst amplitude necessitates a larger detector dynamic range.
-In fact, as mentioned in the principle section, this is a superposition resulting from the convolution of the additional pulse pairs due to low-pass filtering.
-IV.2 Time-domain Signal-to-noise Radio
-The evolution of the time-domain SNR for the H-H-DCS and the C-DCS with respect to the number of averages (or measurement time) is calculated in Fig.
-5(g).
-The x-axis intercepts of the two linear fits exhibit a difference of 2.5575, indicating that this new architecture reduces the required number of averages or measurement time to achieve the same SNR by a factor of 361 (1/361 of the original).
-The y-axis intercepts show a difference of 1.259, indicating an approximate 18-fold enhancement in SNR for the same number of averages.
-These findings are consistent with the expected performance for a setup with m = 20.
-IV.3 Transmission Spectrum
-The enhancement in time-domain SNR is manifested in the transmission spectrum.
+Details regarding the higher-order components are discussed in Supplement 1.
+IV.2 Signal-to-Noise Ratio and Noise Floor
+The power spectral densities (PSDs) of the noise sections of the three DCS are shown in Fig.
+6.
+In a single 1-ms measurement, none of them reached the shot-noise limit.
+This shortfall is primarily due to the optical power loss introduced by bandpass filtering.
+The optical power increased by a factor of about $\sim$17x from C-DCS to PH-PH-DCS, whereas the noise power only increased by a factor of 1.74x (2.4 dB).
+This discrepancy is likely due to quantization noise or relative intensity noise (RIN).
+The evolution of the time-domain SNR for the PH-PH-DCS and the C-DCS as a function of the number of averages (or measurement time) is shown in Fig.
+6 (c).
+The x-axis separation of 2.463 between fit 2 and fit 1 at the same SNR level indicates that this new architecture reduces the averaging time requirement by a factor of 290, i.e., down to 1/290 of the original time.
+The y-axis intercepts differ by 1.222, corresponding to a 16.7-fold SNR enhancement at the same number of averages.
+These observations agree well with the expected performance for a multiplexing factor of m = 20 in approximately constant-noise-limited region.
+Nevertheless, if the system operates in the shot-noise limited region 18, the SNR enhancement and averaging time reduction are expected to decrease to factors of 4 and 17, respectively.
+More generally, for an arbitrary multiplexing factor m, the SNR scales as $\sqrt{m}$ and the averaging time reduces by a factor of m.
 Fig.
-6 presents the acquired transmission spectra around the P7 absorption line of $\mathrm{H^{13}C^{14}N}$ using the three architectures, with a spectral point spacing of 12.5 MHz.
-Both architectures, particularly the H-H-DCS, demonstrate significantly reduced residual levels within the same measurement time compared to the C-DCS.
-Furthermore, the standard deviations of the fitting residual observed in a 5-ms measurement using H-H-DCS are comparable to those in a $\sim$ 2-s measurement using C-DCS.
+6: (a) Noise sections of interferogram (IGM) used to calculate PSDs.
+The main and small centerbursts and their surrounding regions containing the free induction decay are excluded.
+(b) PSDs of the noise sections of the three DCS.
+Three colored dashed lines represent the theoretical shot noise limits for the corresponding DCS in a single measurement (1 ms), calculated from the optical power and detector response.
+The shaded region from 3 to 6 MHz corresponds to the location of the interferogram spectrum.
+Within this region, in a single measurement, the noise power of C-PH-DCS (green) and PH-PH-DCS (red) increased by 0.5 dB and 2.5 dB, respectively, compared to C-DCS (blue).
+In a $\sim$2 s measurement (1999 averages), the increase was 0.75 dB and 2.4 dB, respectively.
+(c) Evolution of time-domain SNR for the PH-PH-DCS and the C-DCS with the number of averages.
+On the log-log scale, the linear fitting models of the two datasets are given by Fit 1 = 0.4851(4)x + 2.475(2) and Fit 2 = 0.4962(4)x + 1.253(1).
+The slopes, approximating 0.5, indicate that the SNR is proportional to the square root of the number of averages, consistent with the principle of coherent averaging.
+The time-domain SNR is defined as the ratio of the centerburst amplitude to the standard deviation of some segments away from the centerburst 44.
+Fig.
+7: Example $\mathrm{H^{13}C^{14}N}$ spectra relative to the center frequency of 193.5449 THz (1548.9555 nm) at a spectral resolution of 12.5 MHz, along with the corresponding weighted Voigt profile fits.
+(a) Recorded in 1.999 s (1999 averages) in the C-DCS.
+(b) Recorded in 1.999 s in the C-PH-DCS.
+(c) Recorded in 1.999 s in the PH-PH-DCS.
+IV.3 Baseline Normalization and Transmission Spectrum
+As mentioned, the baselines of these spectra were normalized by the second reference channel.
+This is because the actual modulated phase levels deviate from the ideal values, leading to mode-to-mode variations in the spectrum, as shown in Figs.
+5 (d) and (f).
+Such variations are also commonly observed in DCS setups based on electro-optic (EO) combs 13 and certain microcombs.
+These high-frequency baseline fluctuations are often difficult to correct using conventional self-fitting methods 45.
+Nevertheless, in PPHE-DCS, such fluctuations exhibit a periodicity of $m$-times mode spacing.
+Utilizing this periodic characteristic, a spectrum of PPHE-DCS can be deinterleaved into m normal spectra, which can then be fitted and processed individually as in C-DCS, without relying on the second channel.
+A detailed demonstration is provided in Supplement 1.
+The enhancement in time-domain SNR is evident in the transmission spectrum.
+Fig.
+7 presents the transmission spectra acquired around the P7 absorption line of $\mathrm{H^{13}C^{14}N}$ with the three architectures, with a spectral point spacing of 12.5 MHz.
+Both architectures, particularly the PH-PH-DCS, demonstrate significantly lower residual levels within the same measurement time compared to the C-DCS, with the residual noise reduced by a factor of $\sim$20.
+It is worth noting that the three weighted Voigt profile fits show slight differences in line depth, we attribute this to the differences in SNR and tracking bandwith of the cross-correlation function 46.
 IV.4 Discussion
-The proof-of-concept results presented in this work can be further improved by combining with ultrahigh-repetition-rate comb sources, or when probing with finer spectral resolution.
-For example, when combined with 22-GHz microresonator combs 45 and operating at a spectral resolution of 500 MHz, compared to conventional DCS with the same resolution, an exceeding 1500-fold averaging-time reduction is expected.
-Furthermore, integrated lithium niobate electro-optic modulators operating at CMOS-compatible voltages may facilitate this implementation 46.
-In wavelength regions where mature EOPMs are not readily available, this approach may can be achieved using nonlinear means.
-For example, under second-harmonic generation 47, a temporal pulse-phase profile of $[0,0,0,\pi/2]$ is anticipated to transform to the desired $[0,0,0,\pi]$, and a profile of $\pi\frac{s}{2m}k^{2}$ is anticipated to transform to the desired $\pi\frac{s}{m}k^{2}$.
-For field-deployed DCS, the mitigation in coherence time inherently offers significant advantages.
-Further possible combination with ultra-low-noise optical frequency combs, such as the compact monolithic solid-state design 48 that provides a $\sim$ 100 Hz linewidth in a quasi-free-running state, complex and sophisticated locking systems may be greatly simplified.
-Moreover, the decoupling of the centerburst repetition rate and the $\Delta f_{\mathrm{rep}}$ (i.e.
-$\Delta f_{\mathrm{s}}$ ) also overcomes the limitation on the noise tracking bandwidth of post-treatment algorithms in free-running DCS, which was previously limited to half the IGM repetition rate (i.e.
-$\Delta f_{\mathrm{rep}}/2$) 43, 49, a restriction that can be roughly attributed to the fact that in conventional DCS, only a single centerburst within each IGM provides effective sampling.
-Finally, it is worth emphasizing that the essence of this spectral mode enhancement lies in enabling the optical frequency comb or combs to output multiple special harmonic pulse trains with the same carrier-envelope offset frequency, but different carrier-envelope phases (non-linear slip), thereby enhancing the majority of the mode sets rather than enhancing a few and suppressing others.
-The scheme demonstrated in this paper, which employs TPM based on EOPMs outside the cavity, may not be the only method to achieve this SNR enhancement.
-For example, in some parametrically driven Kerr soliton microresonators, two soliton states with opposite phases owing to the $Z_{2}$-symmetry have been observed 50, and a soliton sequence with random phase bits of $[0,0,\pi,0]$ has been demonstrated 51.
+We expect this concept can be integrated with ultra-high-repetition-rate OFC source, such as microcombs 47.
+In that case, this PPHE-DCS can achieve a higher SNR enhancement or averaging time reduction, and become applicable to typical use cases requiring spectral sampling spacings from hundreds of MHz to the GHz range 21.
+Finally, it is worth discussing that the PH-OFC share an essential similarity with harmonic mode-locking: the pulse train exhibits correlation between every mth pulse 32, whereas neighboring pulses of which are not fully correlated.
+Under this condition, the spectrum contains an interleaving of m supermodes, resulting in an effective mode spacing of $f_{\mathrm{rep}}/m$ rather than $f_{\mathrm{rep}}$.
+These additional supermodes, apart from the dominant one, are conceptually similar to the ”densified” or ”self-imaging” modes here.
+Therefore, We speculate that the extra-cavity-TPM-modulating scheme, may not be the only viable approach to realize this concept.
+For instance, conventional harmonically mode-locked lasers aim for strong supermode suppression 48.
+In contrast, if low supermode suppression could be achieved such that the m supermodes have comparable intensities, it might provide a suitable source for PPHE-DCS.
+Moreover, in some parametrically driven Kerr soliton microresonators, two soliton states with opposite phases owing to the $Z_{2}$-symmetry have been observed 49, and a soliton sequence with random phase bits of $0,0,\pi,0]$ has been demonstrated 50.
 V Conclusion
-In summary, this work has opened up another effective way to enhance the SNR in DCS beyond coherent averaging, through coherent superposition of harmonic pulses.
-In our proof-of-concept experiments based on temporal phase modulation, we have successfully demonstrated an $\sim$ 18-fold SNR enhancement in a single acquisition, which is equivalent to an SNR gain achieved by coherent averaging over $>$300 periods in conventional DCS.
-This approach avoids the need for long-time averaging, thus enabling recovery of the inherent excellent temporal resolution of DCS and promising to relax stringent requirements on coherence time and stability of DCS systems.
-This could potentially have significant implications for future DCS research and applications.
+In summary, we demonstrate a DCS concept that enhances SNR to reduces coherent averaging time under detector dynamic range limitations by employing phase-pattern higher-repetition-rate frequency combs as sources.
+We reinterpret the self-imaging process of comb spectra from the perspective of mode interference among sub-pulse trains, and provide a comprehensive description of two architectures for this special form of DCS.
+In a proof-of-concept experiment, by 20-fold pulse multiplexing, we achieved an $\sim$16.7-fold SNR improvement or a corresponding $\sim$290-fold reduction in averaging time compared to a conventional DCS architecture, in the non-shot-noise-limited regime.
+These performance gains are expected to decrease to $\sim$4-fold and $\sim$17-fold, respectively, in the shot-noise-limited regime.
+This approach reduces the reliance on long-term averaging, thereby enabling recovery of the inherent temporal resolution of DCS and potentially relaxing the stringent requirements on coherence time and stability of DCS systems.
+We expect that this concept can be combined with ultra-high-repetition-rate microcombs for deployment in applications that typically require spectral sampling spacings from hundreds of MHz to the GHz range.
 Acknowledgments